# Stochastic synaptic dynamics under learning

### v1

@@ Abstract
Learning is based on synaptic plasticity, which drives and is driven by neural activity. Because pre- and postsynaptic spiking activity is shaped by randomness, the synaptic weights follow a stochastic process, requiring a probabilistic framework to capture the noisy synaptic dynamics. We consider a paradigmatic supervised learning example: a presynaptic neural population impinging in a sequence of episodes on a recurrent network of integrate-and-fire neurons through synapses undergoing spike-timing-dependent plasticity (STDP) with additive potentiation and multiplicative depression. We first analytically compute the drift- and diffusion coefficients for a single synapse within a single episode (microscopic dynamics), mapping the true jump process to a Langevin and the associated Fokker-Planck equations. Leveraging new analytical tools, we include spike-time–resolving cross-correlations between pre- and postsynaptic spikes, which corrects substantial deviations seen in standard theories purely based on firing rates. We then apply this microdynamical description to the network setup in which hetero-associations are trained over one-shot episodes into a feed-forward matrix of STDP synapses connecting to neurons of the recurrent network (macroscopic dynamics). By mapping statistically distinct synaptic populations to instances of the single-synapse process above, we self-consistently determine the joint neural and synaptic dynamics and, ultimately, the time course of memory degradation and the memory capacity. We demonstrate that specifically in the relevant case of sparse coding, our theory can quantitatively capture memory capacities which are strongly overestimated if spike-time–resolving cross-correlations are ignored. We conclude with a discussion of the many directions in which our framework can be extended.
@@ I Introduction
Figure 1: We investigate memory properties by studying a single synapse dynamics that captures network properties effectively by a mean-field theory. (a) A Poisson process $\eta$ drives the neuron $x$ through the synapse $w$. The weight $w$ is plastic, i.e., it dynamically depends on $x$ and $\eta$. (b) STDP window Eq. (2) for $w=0.01$ (light gray) through $w=0.2$ (black). The setup (a,b) is studied in Sec. III (c) A cue representation $\boldsymbol{q}$ in a layer of Poisson processes (gray) is activated and drives a recurrent neural network through a matrix of plastic synapses. Among the excitatory neurons (red) in the recurrent network, a target representation $\boldsymbol{p}$ receives an additional stimulus. (d) Training procedure to store hetero-associations between cues and targets. The setup (c,d) is studied in Sec. IV. Parameters $\Delta_{c}=2\times 10^{-3}$, $r_{ac}=8\times 10^{-3}$, $\tau_{c}=16.8/20$, and $\tau_{ac}=33.7/20$ are matched to [5].
@@ II Models
First, we introduce the synaptic model depicted in Fig. 1(a,b). This model depends on pre- and postsynaptic spike trains, thus, next, we define the neuron models used in this paper. Then we introduce the network model depicted in Fig. 1(c), which is composed of the synapse- and neuron models above. We are interested how hetero-associations can be stored into this network; the necessary training scheme [see Fig. 1(d)] is explained last.
@@ II.1 Synapse model
see Fig. 1(b), with $\Delta_{c},r_{ac}>0$ determining the amplitudes and $\tau_{c},\tau_{ac}>0$ being the time scales of potentiation and depression, respectively. In addition to Eq. (1) and Eq. (2), we demand that the considered excitatory synapse maintains a positive weight throughout, $w(t)\geq 0$, a necessary condition for Dale’s law, and implemented in our simulations by a clipping boundary condition. The plasticity rule Eq. (2) with its multiplicative depression and additive potentiation is chosen as it captures the experimental findings of Ref. [5] and because of its inherent stability. Generalizations of Eq. (1), e.g., triplet rules [36, 37] or neuromodulatory dynamics [19] are of interest but out of scope of our study.
@@
This embedding is useful for simulations as it is makes the system local in time. To avoid switch-on effects in Eq. (3), we integrate $A_{\text{pre}}$ and $A_{\text{post}}$ for some time before we start integrating $w$; this way, $A_{\text{pre}}$ and $A_{\text{post}}$ thermalize, i.e., their statistics become stationary, which is in line with the lower integral bound at -$\infty$ in Eq. (1).
@@ II.2 Neuron model
We here introduce a neuron model that is driven (i) by Gaussian white noise and (ii) by a single presynaptic neuron through a single synapse. The dynamics of this synapse will be studied in Sec. III. The dynamics of multiple synapses providing input to a neuron, studied in Sec. IV, can be effectively mapped to the single-synapse case introduced here. The synaptic dynamics Eq. (1) is driven by the pre- and postsynaptic neural activity $\eta$ and $x$, respectively. Thus, to specify the synaptic dynamics, we need to specify the neuron models. Throughout the paper, $\eta(t)$ is the spike train of a Poisson process with rate $\nu$, and $x(t)$ is the spike train of a model neuron driven by $\eta$. The developed framework holds true for arbitrary neuron models $x$, but we give specific expressions and simulation results for the LIF neuron $x(t)=\sum_{i}\delta(t-t_{i}^{\text{post}})$, where $t_{i}^{\text{post}}$ are the times at which the membrane voltage governed by
@@ II.3 Network model
Here, we introduce the network model depicted in Fig. 1(c), which can store hetero-associations, i.e., feed-forward associations between representations in two distinct populations. A layer of $M$ Poisson processes $\eta_{i}$ (gray circles) drives a recurrent network of $N$ LIF neurons [42] (red and blue circles) through a plastic weight matrix $W$. The recurrent network consists of $N_{E}$ excitatory and $N_{I}$ inhibitory neurons, among which only the excitatory neurons are targeted by $W$. Addidionally, each neuron has exactly $C_{E}$ incoming excitatory connections from the recurrent network and $C_{I}$ incoming inhibitory connections with static weights $J_{EE}=J$, $J_{EI}=J_{II}=-gJ$, $J_{IE}=hJ$, parameterized by $J>0$, $g>0$, and $h>0$.
Most input Poisson processes have a low firing rate $\nu_{\text{lo}}$, however, in each training step, a subset of $f_{c}M$ Poisson generators, represented by the binary vector $\boldsymbol{q}\in\left\{0,1\right\}^{M}$ fires with increased rate $\nu_{\text{hi}}$. Simultaneously, a subset of $f_{s}N_{E}$ excitatory neurons, represented by the binary vector $\boldsymbol{p}\in\left\{0,1\right\}^{N_{E}}$, receives direct Poissonian input $\gamma_{i}$ with rate $\nu_{s}$ and static weight $J_{s}$; the other $(1-f_{s})N_{E}$ excitatory neurons do not receive additional input. We are particularly interested in sparse representations, where the activation ratios $f_{c},f_{s}\ll 1$.
@@ III Stochastic synaptic dynamics
The dynamics of a synaptic weight $w$, Eq. (1), is stochastic due to the randomness of $\eta$ and $x$. Trajectories of $w$ can be obtained in simulations, for instance, for the window function Eq. (2), one has to integrate both stochastic equations, Eq. (3) and Eq. (4). The synaptic weight performs a jump process Fig. 2(a1): with each new pre- or post-synaptic spike, a new set of spike pairs is formed that leads to a finite jump, and between jumps the weight is constant. In this section, we develop a simplified description in terms of a Langevin equation and use it to calculate the transient mean and the variance of an ensemble of synapses.
@@ III.1 Drift coefficient
$\mu+\sqrt{2D}\xi(t)+w\eta(t)$
$\approx$ $\mu_{\text{DA}}(\nu)+\sqrt{2D_{\text{DA}}(\nu)}\xi(t),$ (12)
where $\mu_{\text{DA}}(\nu)=\mu+w\nu$ and $D_{\text{DA}}(\nu)=D+w^{2}\nu/2$, the instantaneous firing rate is given by [40]
$r^{-1}=\int_{\frac{v_{r}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}^{\frac{v_{t}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}ds\,e^{s^{2}}[1+\text{erf}(s)],$ (13)
@@ III.2 Diffusion coefficient
Eq. (19) is the infinitesimal diffusion coefficient. For non-infinitesimal $\Delta t$ long enough (specifically, in Fig. 2(c), $\Delta t=10$), the diffusion is captured by the Langevin dynamics based on Eq. (17) and Eq. (19), as we derive in the next section.
@@ III.3 Mid- and long-term evolution of an ensemble of synapses
The drift- and diffusion coefficients derived above describe the process on infinitesimal times. Here, we study the mean and variance of an ensemble of synapses on finite times. We provide a technically detailed analysis because the results of this section are the main building blocks of the mean-field theory of learning discussed in Sec. IV. Although the dynamics can admit a stationary solution [see e.g. the zero of $D^{(1)}(w)$, i.e., the black solid line crosses zero in Fig. 2(b3-b4)], we are here mainly interested in transient dynamics: If the training patterns permanently (but slowly) change and the synapse is subject to slow homeostatic plasticity, the stationary state will never be reached in a learning situation. Therefore we here study the non-equilibrium dynamics of the ensemble mean and variance.
@@
Eq. (24) and Eq. (26) can be efficiently evaluated to obtain the mean and variance of an ensemble of synapses. These equations are thus useful formulations of the microscopic (i.e., single-synapse) model, which we leverage in the next section to derive macroscopic (i.e., population) dynamics. We also use Eq. (24) and Eq. (26) to verify the second moment of the stochastic increment over a finite time window Fig. 2(c); i.e., due to the invalidity of the infinitesimal diffusion coefficient, we predict the finite-time diffusion coefficient as put forward in [50].
@@ IV Dynamics of learning
In this section we study the network and learning scheme sketched in Fig. 1(c,d) and introduced in Sec. II.3 and Sec. II.4 in Sec. II. The dynamics can now be viewed on two time scales: First, on the discrete macroscopic time scale (pattern index $k$, going from one training session to the next), the data that the network is exposed to is exchanged. Since subsequent pattern pairs are uncorrelated, the sequence of matrices $W_{k}$ after the $k$th training session including homeostasis is a Markov chain (see also [51]). Second, on the continuous microscopic time scale (within a training session), the present pattern pair $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$ appears static, and the fast neural and synaptic dynamics evolve as defined in Sec. II. As we discuss in detail in this section, the system becomes stationary on the macroscopic time scale while remaining non-equilibrium on the microscopic time scale. The following analysis starts by studying the storage of a memory, namely the association $\boldsymbol{q}_{k}\rightarrow\boldsymbol{p}_{k}$. This storage occurs on the microscopic time scale by exposing the system to $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$. The analysis proceeds by quantifying the stationary regime on the macroscopic time scale. Assuming this regime, we lastly study how memories are forgotten on the macroscopic time scale.
@@ IV.1 Synaptic dynamics on the microscopic time
As one can see from the total effective input Eq. (30), the postsynaptic neurons are, within the employed approximation, effectively decoupled apart from their common dependence on the mean fields $m_{ab}$, $V_{ab}$, $r_{E}$, and $r_{I}$, which we now determine self-consistently. Thus, assuming knowledge of the numerical values of the mean fields, the single-neuron statistics is readily determined and we can proceed as in Sec. III: The drift- and diffusion coefficients of the four distinct synaptic populations are, respectively
@@
Lastly, we may close the self-consistency by identifying $m_{ab}$ and $V_{ab}$ with the ensemble average of synapses with drift- and diffusion coefficients Eq. (31). Employing the method from Sec. III.3 in Sec. III, this leads to
@@ IV.3 Forgetting on the macroscopic time scale
Figure 4: Dynamics of memory traces and memory capacity. (a) Mean synaptic strength from cue to (non) target neurons after $k$ subsequently stored associations. Simulation results denoted by circles, green for cue-to-target and red for cue-to-non-target, dark colors are averages over 100100 realizations, light colors are single realizations. Theory Eq. (45) (black lines). (b) Variance of synaptic strength from cue to (non) target. Colors as in (a), theory (black lines) from Eq. (46) and the horizontal line is the stationary variance Eq. (39). (c) Fraction of false positives on average over 100 realizations (dark purple circles), for a single realization (light circles) and theory Eq. (53) based on the full drift (black solid line) and neglecting cross-correlations (dotted line). (d,e) Memory capacity Eq. (48) as a function of input sparseness $f_{c}$ (d) and the input width $M$ (e) with (solid) and without (dotted) cross-correlations. Parameters as in Fig. 3, but $\nu_{\text{lo}}=0$.
@@
Trace degradation
We assume that the system is the macroscopically stationary state identified above. After the entraining of a specific pattern pair, say $(\boldsymbol{q}_{0},\boldsymbol{p}_{0})$, the weights from $\boldsymbol{q}_{0}$ to $\boldsymbol{p}_{0}$ have mean $K_{11}^{(0)}=\gamma_{1}m_{11}(T)$ and variance $G_{11}^{(0)}=\gamma_{1}^{2}V_{11}(T)$. Analogously, the four distinct populations have mean and variance

### v2

@@ Abstract
Learning is based on synaptic plasticity, which affects and is driven by neural activity. Because pre- and postsynaptic spiking activity is shaped by randomness, the synaptic weights follow a stochastic process, requiring a probabilistic framework to capture the noisy synaptic dynamics. We consider a paradigmatic supervised learning example: a presynaptic neural population impinging in a sequence of episodes on a recurrent network of integrate-and-fire neurons through synapses undergoing spike-timing-dependent plasticity (STDP) with additive potentiation and multiplicative depression. We first analytically compute the drift- and diffusion coefficients for a single synapse within a single episode (microscopic dynamics), mapping the true jump process to a Langevin and the associated Fokker-Planck equations. Leveraging new analytical tools, we include spike-time–resolving cross-correlations between pre- and postsynaptic spikes, which corrects substantial deviations seen in standard theories purely based on firing rates. We then apply this microdynamical description to the network setup in which hetero-associations are trained over one-shot episodes into a feed-forward matrix of STDP synapses connecting to neurons of the recurrent network (macroscopic dynamics). By mapping statistically distinct synaptic populations to instances of the single-synapse process above, we self-consistently determine the joint neural and synaptic dynamics and, ultimately, the time course of memory degradation and the memory capacity. We demonstrate that specifically in the relevant case of sparse coding, our theory can quantitatively capture memory capacities which are strongly overestimated if spike-time–resolving cross-correlations are ignored. We conclude with a discussion of the many directions in which our framework can be extended.
@@ I Introduction
Figure 1: We investigate memory properties by studying a single synapse dynamics that effectively captures synaptic population dynamics. (a) A Poisson process $\eta$ drives the neuron $x$ through the synapse $w$. The weight $w$ is plastic, i.e., it dynamically depends on $x$ and $\eta$. (b) STDP window Eq. (2) for $w=0.01$ (light gray) through $w=0.2$ (black). The setup (a,b) is studied in Sec. III (c) A cue representation $\boldsymbol{q}$ in a layer of Poisson processes (gray) is activated and drives a recurrent neural network through a matrix of plastic synapses. Among the excitatory neurons (red) in the recurrent network, a target representation $\boldsymbol{p}$ receives an additional stimulus. (d) Training procedure to store hetero-associations between cues and targets. The setup (c,d) is studied in Sec. IV. Parameters $\Delta_{c}=2\times 10^{-3}$, $r_{ac}=8\times 10^{-3}$, $\tau_{c}=16.8/20$, and $\tau_{ac}=33.7/20$ are matched to [5].
@@ II Models
First, we introduce the synaptic model depicted in Fig. 1(a,b). This model depends on pre- and postsynaptic spike trains, thus, next, we define the neuron models used in this paper. Then we introduce the network model depicted in Fig. 1(c), which is composed of the synapse- and neuron models above. We are interested in how hetero-associations can be stored into this network; the necessary training scheme [see Fig. 1(d)] is explained last.
@@ II.1 Synapse model
see Fig. 1(b), with $\Delta_{c},r_{ac}>0$ determining the amplitudes and $\tau_{c},\tau_{ac}>0$ being the time scales of potentiation and depression, respectively. In addition to Eqs. (1) and (2), we demand that the considered excitatory synapse maintains a positive weight throughout, $w(t)\geq 0$, a necessary condition for Dale’s law, and implemented in our simulations by a clipping boundary condition. The plasticity rule Eq. (2) with its multiplicative depression and additive potentiation is chosen as it captures the experimental findings of Ref. [5] and because of its inherent stability. Generalizations of Eq. (1), e.g., triplet rules [36, 37] or neuromodulatory dynamics [19] are of interest but out of scope of our study.
@@
This embedding is useful for simulations as it makes the system local in time. To avoid switch-on effects in Eq. (3), we integrate $A_{\text{pre}}$ and $A_{\text{post}}$ for some time before we start integrating $w$; this way, $A_{\text{pre}}$ and $A_{\text{post}}$ thermalize, i.e., their statistics become stationary, which is in line with the lower integral bound at -$\infty$ in Eq. (1).
@@ II.2 Neuron model
We here introduce a neuron model that is driven (i) by Gaussian white noise and (ii) by a single presynaptic neuron through a single synapse. The dynamics of this single synapse will be studied in Sec. III. The dynamics of multiple synapses providing input to a neuron, studied in Sec. IV, can be effectively mapped to the single-synapse case introduced here. The synaptic dynamics Eq. (1) is driven by the pre- and postsynaptic neural activity $\eta$ and $x$, respectively. Thus, to specify the synaptic dynamics, we need to specify the neuron models. Throughout the paper, $\eta(t)$ is the spike train of a Poisson process with rate $\nu$, and $x(t)$ is the spike train of a model neuron driven by $\eta$. The developed framework holds true for arbitrary neuron models $x$, but we give specific expressions and simulation results for the LIF neuron $x(t)=\sum_{i}\delta(t-t_{i}^{\text{post}})$, where $t_{i}^{\text{post}}$ are the times at which the membrane voltage governed by
@@ II.3 Network model
Here, we introduce the network model depicted in Fig. 1(c), which can store hetero-associations, i.e., feed-forward associations between representations in two distinct populations. A layer of $M$ Poisson processes $\eta_{i}$ (gray circles) drives a recurrent network of $N$ LIF neurons [42] (red and blue circles) through a plastic weight matrix $W$. The recurrent network consists of $N_{E}$ excitatory (red) and $N_{I}$ inhibitory (blue) neurons, among which only the excitatory neurons are targeted by $W$. Addidionally, each neuron has exactly $C_{E}$ incoming excitatory connections from the recurrent network and $C_{I}$ incoming inhibitory connections with static weights $J_{EE}=J$, $J_{EI}=J_{II}=-gJ$, $J_{IE}=hJ$, parameterized by $J>0$, $g>0$, and $h>0$.
Most input Poisson processes have a low firing rate $\nu_{\text{lo}}$, however, in each training step, a subset of $f_{c}M$ Poisson generators, represented by the ones in binary vector $\boldsymbol{q}\in\left\{0,1\right\}^{M}$ fires with increased rate $\nu_{\text{hi}}$. Simultaneously, a subset of $f_{s}N_{E}$ excitatory neurons, represented by the binary vector $\boldsymbol{p}\in\left\{0,1\right\}^{N_{E}}$, receives direct Poissonian input $\gamma_{i}$ with rate $\nu_{s}$ and static weight $J_{s}$; the other $(1-f_{s})N_{E}$ excitatory neurons do not receive additional input. We are particularly interested in sparse representations, where the activation ratios $f_{c},f_{s}\ll 1$.
@@ III Stochastic synaptic dynamics
The dynamics of a synaptic weight $w$, Eq. (1), is stochastic due to the randomness of $\eta$ and $x$. Trajectories of $w$ can be obtained in simulations, for instance, for the window function Eq. (2), one has to integrate both stochastic equations, Eqs. (3) and (4). The synaptic weight performs a jump process Fig. 2(a1): with each new pre- or post-synaptic spike, a new set of spike pairs is formed that leads to a finite jump, and between jumps the weight is constant. In this section, we develop a simplified description in terms of a Langevin equation and use it to calculate the transient mean and the variance of an ensemble of synapses.
@@ III.1 Drift coefficient
$\mu+\sqrt{2D}\xi(t)+w\eta(t)\approx\mu_{\text{DA}}+\sqrt{2D_{\text{DA}}}\xi(t),$ (12)
where $\mu_{\text{DA}}=\mu+w\nu$ and $D_{\text{DA}}=D+w^{2}\nu/2$, the instantaneous firing rate is given by [40]
$\frac{1}{r(w)}=\int_{\frac{v_{r}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}^{\frac{v_{t}-\mu_{\text{DA}}}{\sqrt{2D_{\text{DA}}}}}ds\,e^{s^{2}}[1+\text{erf}(s)],$ (13)
@@ III.2 Diffusion coefficient
Eq. (19) is the infinitesimal diffusion coefficient. For non-infinitesimal $\Delta t$ long enough (specifically, in Fig. 2(c), $\Delta t=10$), the diffusion is captured by the Langevin dynamics based on Eqs. (17) and (19), as we derive in the next section.
@@ III.3 Mid- and long-term evolution of an ensemble of synapses
The drift- and diffusion coefficients derived above describe the process on infinitesimal times. Here, we study the mean and variance of an ensemble of synapses on finite times. We provide a technically detailed analysis because the results of this section are the main building blocks of the mean-field theory of learning discussed in the next section (Sec. IV). Although the dynamics can admit a stationary solution [see e.g. the zero of $D^{(1)}(w)$, i.e., the black solid line crosses zero in Fig. 2(b3-b4)], we are here mainly interested in transient dynamics: If the training patterns permanently (but slowly) change and the synapse is subject to slow homeostatic plasticity, the stationary state will never be reached in a learning situation. Therefore we here study the non-equilibrium dynamics of the ensemble mean and variance.
@@
Equations (24) and (26) can be efficiently evaluated to obtain the mean and variance of an ensemble of synapses. These equations are thus useful formulations of the microscopic (i.e., single-synapse) model, which we leverage in the next section to derive macroscopic (i.e., population) dynamics. We also use Eqs. (24) and (26) to verify the second moment of the stochastic increment over a finite time window Fig. 2(c); i.e., due to the invalidity of the infinitesimal diffusion coefficient, we predict the finite-time diffusion coefficient as put forward in [50].
@@ IV Dynamics of learning
In this section we study the network and learning scheme sketched in Fig. 1(c,d) (introduced in Secs. II.3 and II.4). The dynamics can now be viewed on two time scales: First, on the discrete macroscopic time scale (pattern index $k$, going from one training session to the next), the data that the network is exposed to is exchanged. Since subsequent pattern pairs are uncorrelated, the sequence of matrices $W_{k}$ after the $k$th training session including homeostasis is a Markov chain (see also [51]). Second, on the continuous microscopic time scale (within a training session), the present pattern pair $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$ appears static, and the fast neural and synaptic dynamics evolve as defined in Sec. II. As we discuss in detail in this section, the system becomes stationary on the macroscopic time scale while remaining non-equilibrium on the microscopic time scale. The following analysis starts by studying the storage of a memory, namely the association $\boldsymbol{q}_{k}\rightarrow\boldsymbol{p}_{k}$. This storage occurs on the microscopic time scale by exposing the system to $(\boldsymbol{q}_{k},\boldsymbol{p}_{k})$. The analysis proceeds by quantifying the stationary regime on the macroscopic time scale. Assuming this regime, we lastly study how memories are forgotten on the macroscopic time scale.
@@ IV.1 Synaptic dynamics on the microscopic time
As one can see from the total effective input Eq. (30), the postsynaptic neurons are, within the employed approximation, effectively decoupled apart from their common dependence on the mean fields $m_{ab}$, $V_{ab}$, $r_{E}$, and $r_{I}$, which we now determine self-consistently. Thus, assuming knowledge of the numerical values of the mean fields, the single-neuron statistics is readily determined and we can proceed as for the single synapse in Sec. III: The drift- and diffusion coefficients of the four distinct synaptic populations are, respectively
@@
Lastly, we may close the self-consistency by identifying $m_{ab}$ and $V_{ab}$ with the ensemble average of synapses with drift- and diffusion coefficients Eq. (31). Employing the method from Sec. III.3, this leads to
@@ IV.3 Forgetting on the macroscopic time scale
Figure 4: Dynamics of memory traces and memory capacity. (a) Mean synaptic strength from cue to (non-)target neurons after $k$ subsequently stored associations. Simulation results denoted by circles, green for cue-to-target and red for cue-to-non-target, dark colors are averages over 100100 realizations, light colors are single realizations. Theory Eq. (45) (black lines). (b) Variance of synaptic strength from cue to (non-)target. Colors as in (a), theory (black lines) from Eq. (46) and the horizontal line is the stationary variance Eq. (39). (c) Fraction of false positives on average over 100 realizations (dark purple circles), for a single realization (light circles) and theory Eq. (53) based on the full drift (black solid line) and neglecting cross-correlations (dotted line). (d,e) Memory capacity Eq. (48) as a function of input sparseness $f_{c}$ (d) and the input width $M$ (e) with (solid) and without (dotted) cross-correlations. Parameters as in Fig. 3, but $\nu_{\text{lo}}=0$.
@@
IV.3.1 Trace degradation
We assume that the system is in the macroscopically stationary state identified above. After the entraining of a specific pattern pair, say $(\boldsymbol{q}_{0},\boldsymbol{p}_{0})$, the weights from $\boldsymbol{q}_{0}$ to $\boldsymbol{p}_{0}$ have mean $K_{11}^{(0)}=\gamma_{1}m_{11}(T)$ and variance $G_{11}^{(0)}=\gamma_{1}^{2}V_{11}(T)$. Analogously, the four distinct populations have mean and variance